OpenStudy (anonymous):

7(2x + 6) = 1 is it -41/14?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Distribute: 7*(2x) + 7*6 = 1 Simplify: 14x + 42 = 1 Subtract 42 from each side of the equation: 14x = -41 Divide by 14: x = -41/14 You are correct.